Lithium atom has two naturally occuring isotopes 26Li and 27Li. Given that the relative molecular mass of Lithium is 6.94. What is the abundance of these isotopes?


1
Expert's answer
2021-08-12T06:44:44-0400

average atomic mass = (26 + 27)/2

= 26.5

26.5% = x + (1-x)

26.5 × 3.47 = 91.955

91.955 = 2x

X = 45.978